OpenStudy (anonymous):

please help me with problem anyone! Solve x2 + 4x + 8 = 0.

OpenStudy (mrnood):

in standard form a quadratic is written ax^2+bx+c=0 You may know the quadratic formula. However if the 'discriminant' of the equation is negative then there is no real solution. The discriminant = b^2 - 4ac Are you studying complex numbers? if so - the result is complex - if not - there is no real solution

OpenStudy (anonymous):

yes i'm studying complex numbers

OpenStudy (mrnood):

OK - then write the quadratic formula to get the solutions You will end up with sqrt (-n) in the formula - so simplify that using sqrt(-n) = i sqrt(n)

OpenStudy (anonymous):

i got \[-4 \pm \sqrt{-16}\div2\]

OpenStudy (anonymous):

x equals negative 2 plus or minus 4 I is this the answer

OpenStudy (mrnood):

not quite - you need to divide the WHOLE by 2

OpenStudy (anonymous):

x equals negative 2 plus or minus 4 I is this right?

OpenStudy (mrnood):

\[x= \frac{-4\pm \sqrt{-16} }{2 }\] not quite - you need to divide the WHOLE by 2

OpenStudy (anonymous):

my answer choices are x equals negative 2 plus or minus 2 I x equals negative 2 plus or minus 4 I x equals negative 4 plus or minus 2 I x equals negative 4 plus or minus 4 I

OpenStudy (mrnood):

so - you are nearly there - youtr first answer was wrong for the reason I told you. Have another look at your arithmetic

OpenStudy (anonymous):

that's confusing what

OpenStudy (mrnood):

all I am saying is that you have done all the hard work - but you made a mistake with the division Have another look at my formula and simplify it

OpenStudy (anonymous):

would it be x equals negative 4 plus or minus 2 I

OpenStudy (mrnood):

have confidence - is that correct?

OpenStudy (anonymous):

I don't know...

OpenStudy (mrnood):

\[\frac{ -4 }{ 2 }\pm \frac{ \sqrt{16} }{ 2 }\]

OpenStudy (mrnood):

sorry - should be -16 above....

OpenStudy (mrnood):

you have done all the work - and your answer IS correct. But - you obviously understand quadratics and complex nos - so you should decide for yourself when you have the right answer...

OpenStudy (anonymous):

so of the following answer choices which one is correct x equals negative 2 plus or minus 2 I x equals negative 2 plus or minus 4 I x equals negative 4 plus or minus 2 I x equals negative 4 plus or minus 4 I

OpenStudy (mrnood):

you have done all the work - and your answer IS correct. But - you obviously understand quadratics and complex nos - so you should decide for yourself when you have the right answer...